# Microwave Output Stabilization of a Qubit Controller via Device-Level Temperature Control

### v1

@@ Abstract
We present the design and performance of QuEL-1 SE, which is a multichannel qubit controller developed for superconducting qubits. The system incorporates the active thermal stabilization of critical analog integrated circuits, such as phase-locked loops, amplifiers, and mixers, to suppress the long-term amplitude and phase drift. To evaluate the amplitude and phase stability, we simultaneously monitor 15 microwave output channels over 24 h using a common analog-to-digital converter. Across the channels, the normalized amplitude exhibits standard deviations of 0.09%–0.22% (mean: 0.15%), and the phase deviations are 0.35∘–0.44∘ (mean: 0.39∘). We further assess the impact of these deviations on quantum gate operations by estimating the average fidelity of an gate under the coherent errors corresponding to the deviations. The resulting gate infidelities are for amplitude errors and for phase errors, which are significantly lower than typical fault-tolerance thresholds such as those of the surface code. These results demonstrate that the amplitude and phase stability of QuEL-1 SE enables reliable long-duration quantum operations, thus highlighting its utility as a scalable control platform for superconducting and other qubit modalities.
@@ I Introduction
These works have emphasized the high sampling rate of digital-to-analog converters (DACs) and analog-to-digital converters (ADCs), low spurious-free dynamic range, and low phase noise [27, 28]. However, negligible attention has been paid to the long-term stability of generated microwave signals. As the number of qubits increases and quantum error correction becomes feasible, which enable long-duration quantum operations, the stability of the microwave amplitude and relative phase becomes critical for achieving high-fidelity quantum gates. For example, although QICK [18] shows phase stability, it is demonstrated by resetting the phase for each short sequence. This limits its applicability to long sequences and systems, including analog up/downconversions. Similarly, in M2CS [21], the amplitude stability is reported only for a single output signal. In addition, the sources and mitigation strategies for long-term variations in the microwave amplitude and phase have not been discussed in detail.
@@ II.1 Clocking and Frequency Reference Infrastructure
For this purpose, we develop a clock distributor that provides three types of clock signals to twelve channels, resulting in thirty-six output ports. Two distributor configurations are implemented: a primary distributor, which uses a 10 MHz rubidium oscillator as its reference, and a secondary distributor, which derives its input from one set of the three distributed clocks. This architecture enables the scalable expansion of the QuEL-1 SE systems. Figure 2 illustrates the overall clock distribution network and its connection to the QuEL-1 SE units.
@@ III Temperature Stabilization and Active Thermal Management
The temperature of each thermistor is digitized and used as an input to a proportional–integral compensator, which is individually implemented on a processor. The result of the compensator is used as the input of a pulse width modulation (PWM) amplifier to drive an actuator that is either a fan or a heater. The block diagram of the control system is shown in Fig. 4 This hybrid approach minimizes system-wide and component-level thermal variations, thereby contributing to the overall performance and long-term reliability of the controller.
@@ IV.1 Experimental Configuration
There are two reasons for selecting only CTRL channels and not using the ROUT channels. First, an ROUT channel is always used with an RIN channel that shares a common PLL IC. Thus, the microwave fluctuations of the ROUT channels are consistently smaller than those of the CTRL channels. Second, although the signals from the ROUT channels pass through the upconversion stage, which is not included in the CTRL channels, the temperature variations in the mixer and LO (LMX2594) are already considered in the CTRL channels in the downconversion stage before digitization.
@@ V Discussion and Conclusion
Using Equation (1) with and gives
where the last expression is the small-angle approximation. With the measured standard deviation, ,
@@
Equation (1) yields
where the last expression holds for small values of . With the measured standard deviation, ,
In summary, at the level of the observed fluctuations, the -gate infidelity due to the amplitude noise () and phase misalignment () is considerably small. These values are significantly lower than the typical fault-tolerance targets for single-qubit operations, indicating that the measured amplitude and phase stability have a negligible impact on the -gate performance.

### v2

@@ Abstract
We present the design and performance of QuEL-1 SE, which is a multichannel qubit controller developed for superconducting qubits. The system incorporates the active thermal stabilization of critical analog integrated circuits, such as phase-locked loops, amplifiers, and mixers, to suppress the long-term amplitude and phase drift. To evaluate the amplitude and phase stability, we simultaneously monitor 15 microwave output channels over 24 h using a common analog-to-digital converter. Across the channels, the normalized amplitude exhibits standard deviations of 0.09%–0.22% (mean: 0.15%), and the phase deviations are 0.35∘–0.44∘ (mean: 0.39∘). We further assess the impact of these deviations on quantum gate operations by estimating the average fidelity of an gate under the coherent errors corresponding to the deviations. The resulting estimated contribution to the gate infidelity due to amplitude noise () and phase mis-alignment () is significantly lower than typical fault-tolerance thresholds such as those of the surface[Fowler_2012]. These results demonstrate that the amplitude and phase stability of QuEL-1 SE enables reliable long-duration quantum operations, thus highlighting its utility as a scalable control platform for superconducting and other qubit modalities.
@@ I Introduction
These works have emphasized the high sampling rate of digital-to-analog converters (DACs) and analog-to-digital converters (ADCs), low spurious-free dynamic range, and low phase noise [27, 23]. However, negligible attention has been paid to the long-term stability of generated microwave signals. As the number of qubits increases and quantum error correction becomes feasible, which enable long-duration quantum operations, the stability of the microwave amplitude and relative phase becomes critical for achieving high-fidelity quantum gates.
For example, the QICK [7] control platform employs direct digital synthesis with phase reset at each experimental repetition to deterministically define the initial phase of all drives. This functionality ensures reproducible phase relationships within individual experiments. In contrast, the problem considered here concerns long-timescale operation, where phase coherence must persist over durations relevant to repeated error-correction cycles and system-level operation. Similarly, in M2CS [37], the amplitude stability is reported only for a single output signal. In addition, the sources and mitigation strategies for long-term variations in the microwave amplitude and phase have not been discussed in detail.
@@ II.1 Clocking and Frequency Reference Infrastructure
For this purpose, we develop a clock distributor that provides three types of clock signals to twelve channels, resulting in thirty-six output ports. Two cascaded distributors are implemented: a primary distributor, which uses a 10 MHz rubidium oscillator as its reference, and a secondary distributor, which derives its input from one set of the three distributed clocks. This architecture enables the scalable expansion of the QuEL-1 SE systems. Figure 2 illustrates the overall clock distribution network and its connection to the QuEL-1 SE units.
@@ III Temperature Stabilization and Active Thermal Management
The temperature of each thermistor is digitized and used as an input to a proportional–integral compensator, which is individually implemented in a processor. The result of the compensator is used as the input of a pulse width modulation (PWM) amplifier to drive an actuator that is either a fan or a heater. The block diagram of the control system is shown in Fig. 4 This hybrid approach minimizes system-wide and component-level thermal variations, thereby contributing to the overall performance and long-term reliability of the controller.
@@ IV.1 Experimental Configuration
There are two reasons for selecting only CTRL channels and not using the ROUT channels. First, an ROUT channel is always used with an RIN channel that shares a common PLL IC. Thus, measuring ROUT with RIN underestimates the actual microwave fluctuations due to correlated PLL noise, leading to an unfair comparison with CTRL channels. Second, although the signals from the ROUT channels pass through the upconversion stage, which is not included in the CTRL channels, the temperature variations in the mixer and LO (LMX2594) are already considered in the CTRL channels in the downconversion stage before digitization as shown in Fig. 1.
@@ V Discussion and Conclusion
[…]
where the last expression is the small-angle approximation. With the measured standard deviation, , , whereas without thermal control () it increases to .
@@
[…]
where the last expression holds for small values of . With the measured standard deviation, , , whereas without thermal control () it increases to .
In summary, based on the observed amplitude and phase fluctuations, the estimated contribution to the -gate infidelity due to amplitude noise () and phase misalignment () is small. While these values are not experimentally verified through direct qubit measurements, they suggest that the achieved microwave stability would not be a limiting factor for single-qubit gate performance, even for future superconducting qubits with substantially improved coherence times.